# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v1

@@ 3.2 Contraction and mixing
$\mu_{h-\epsilon}^{\leftarrow}(x)=\tilde{\mu}_{h-\epsilon}^{\leftarrow}(x)=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(x)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
This follows from that if we consider system started at time $\epsilon>0$, with initial distribution $\mu_{\epsilon}$, then we obtain the above through the Bayesian perspective on the forward and reverse heat semigroups, elaborated in Appendix D.
We now show that the following integral is indeed integrable, so $\tilde{\mu}_{h}^{\leftarrow}$ is well-defined:
$\tilde{\mu}_{h}^{\leftarrow}(x):=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}\nu(x)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
For fixed $x$ and $\epsilon<h/2$,
$\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu(z)\,\mathrm{d}z\gtrsim\exp\bigl{(}-\frac{(\|y-x_{0}\|+D)^{2}}{2(h-\epsilon)}\bigr{)}\,,$
as the support of $\nu$ is constrained to $\mathcal{K}\subset B_{D}(x_{0})$. Since $\mu_{0}$ has bounded support, $\mu_{h}(y)\lesssim\exp(-\frac{\|y\|^{2}}{a})$ for some constant $a>0$. Thus,
$\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\,\mu_{h}(y)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\lesssim\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\,\mu_{h}(y)}{\exp(-\frac{(\|y-x_{0}\|+D)^{2}}{2(h-\epsilon)})}\lesssim\exp\Bigl{(}\frac{\langle 2(x-x_{0}),y\rangle+2D\|y-x_{0}\|}{h}-\frac{\|y\|^{2}}{a}\Bigr{)}\,,$
and the last bound is integrable in $y$.
We then show the pointwise convergence of $\tilde{\mu}_{h-\epsilon}^{\leftarrow}$ to $\tilde{\mu}_{h}^{\leftarrow}$ as $\epsilon\to 0$. Note that $\nu P_{\epsilon}\to\nu$, as $\nu$ has a bounded support. Also, the denominator is independent of $\epsilon$ due to
$\frac{1}{(2\pi(h-\epsilon))^{d/2}}\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z=\mathcal{N}(0,(h-\epsilon)I)*\nu P_{\epsilon}=\nu*\mathcal{N}(0,hI)\,.$
@@
$\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\leq\Bigl{(}\frac{h}{h-\epsilon}\Bigr{)}^{d/2}\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\lesssim\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,.$
As shown above, the last bound is integrable with respect to $\mu_{h}$, so the dominated convergence theorem implies
$\lim_{\epsilon\to 0}\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,,$
Thus, the pointwise convergence follows. Note that if we take $\nu(x)=\pi^{X}(x)=\frac{\mathds{1}_{\mathcal{K}}(x)}{\operatorname{vol}(\mathcal{K})}$, then $\tilde{\mu}_{h}^{\leftarrow}$ is the distribution of the backwards step of our algorithm. In particular, this corresponds to first sampling $x\sim\mu_{h}$, then $y\sim Q_{h}^{\nu,h}(\cdot|x)$, which is precisely the law of $\mu_{h}^{\leftarrow}$ given by ($\mathsf{FP\text{-}BH}$).
As for the second statement, it follows from Scheffé’s lemma [billingsley2017probability, Theorem 16.12] that the pointwise convergence of $\mu_{h-\varepsilon}^{\leftarrow}\to\mu_{h}^{\leftarrow}$ leads to its $\mathsf{TV}$-convergence, which in turn implies the weak convergence. It follows from lower semicontinuity of $D_{f}$ [ambrosio2000functions, Theorem 2.34] that the weak convergence ensures $D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu^{\leftarrow}_{h-t})$. ∎
@@ 3.3 Failure probability and wasted steps
We begin by defining a suitable version of local conductance [kannan1997random].
@@
$\ell(x)\stackrel{{\scriptstyle\mathrm{{\scriptscriptstyle def}}}}{{=}}\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}{\int_{\mathbb{R}^{d}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}=\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}{(2\pi h)^{d/2}}\,.$
@@
$\pi^{Y}(\mathcal{K}_{\delta}^{c})\leq e^{c_{1}^{2}}\exp(-\nicefrac{{t^{2}}}{{8c_{1}^{2}}})$ for $\delta=t/d$ and variance $h=c_{1}^{2}/d^{2}$, where $c_{1}$ is some constant and $t\geq 2c_{1}(c_{1}+1)$.
@@
For $y\in\partial\mathcal{K}_{\delta}$, we can take the supporting half-space $H$ at $\mathsf{proj}_{\mathcal{K}}(y)$ containing $\mathcal{K}$, due to convexity of $\mathcal{K}$. Then,
$\pi^{Y}(\mathcal{K}_{\delta}^{c})$ $=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{\mathcal{K}}\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y\leq\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{H}\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y$
$=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y\,.$ (3.3)
@@
$\mathsf{T}(s):=\mathbb{P}_{Z\sim\mathcal{N}(0,h)}(Z\geq s)=\frac{1}{\sqrt{2\pi h}}\int_{s}^{\infty}\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}\,\mathrm{d}z\,.$
By the co-area formula and integration by parts,
$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\mathrm{d}y=\int_{\delta}^{\infty}\mathsf{T}(s)\operatorname{vol}(\partial\mathcal{K}_{s})\,\mathrm{d}s$
@@
Recall that $\mathsf{T}(s)\leq\frac{1}{2}\exp(-\frac{1}{2}(sh^{-1/2})^{2})$ for $s\geq\delta=t/d$ due to a standard tail bound on a Gaussian distribution. This tail bound, combined with
@@
$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y$ $\leq\frac{1}{\sqrt{2\pi h}}\int_{\delta}^{\infty}\exp\bigl{(}-\frac{s^{2}}{2h}\bigr{)}\bigl{(}\underbrace{(1+s)^{d}}_{\leq\exp(sd)}-1\bigr{)}\operatorname{vol}(\mathcal{K})\,\mathrm{d}s$
@@
$\underset{(i)}{\leq}\operatorname{vol}(\mathcal{K})\exp(hd^{2}/2)\exp\bigl{(}-\frac{1}{2}\,(\delta h^{-1/2}-dh^{1/2})^{2}\bigr{)}$
$\underset{(ii)}{\leq}\operatorname{vol}(\mathcal{K})\exp(c_{1}^{2}/2)\exp\bigl{(}-\frac{t^{2}}{8c_{1}^{2}}\bigr{)}\,,$
where in $(i)$ we used the tail bound of a Gaussian, and $(ii)$ follows from that in the regime of $h=c_{1}^{2}d^{-2}$, we have $\delta h^{-1/2}-dh^{1/2}-1\geq t/c_{1}-(c_{1}+1)\geq\nicefrac{{t}}{{2c_{1}}}$ for $t\geq 2c_{1}(c_{1}+1)$. Putting the last bound into (3.3) completes the proof of the claim. ∎
@@
$\ell(y)\leq\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z=\mathbb{P}_{Z\sim\mathcal{N}(0,h)}(Z\geq\delta)\leq\exp(-\Omega(t^{2}))\,.$
@@
$\leq\pi^{Y}(\mathcal{K}_{\delta}^{c})+\int_{[\ell\geq N^{-1}\log(3mM/\eta)]}\exp(-\ell N)\,\mathrm{d}\pi^{Y}+\int_{\mathcal{K}_{\delta}\cap[\ell<N^{-1}\log(3mM/\eta)]}\frac{\ell(y)}{\operatorname{vol}(\mathcal{K})}\,\mathrm{d}y$
$\leq e^{c_{1}^{2}}\exp(-\frac{t^{2}}{8c_{1}^{2}})+\frac{\eta}{3mM}+\frac{\log(3mM/\eta)}{N}\,\frac{\operatorname{vol}(\mathcal{K}_{\delta})}{\operatorname{vol}(\mathcal{K})}$
$\leq e^{c_{1}^{2}}\exp(-\frac{t^{2}}{8c_{1}^{2}})+\frac{\eta}{3mM}+\frac{e^{t}}{N}\,\log\frac{3mM}{\eta}\,,$
where we used $\operatorname{vol}(\mathcal{K}_{\delta})\subset\operatorname{vol}\bigl{(}(1+\delta)\mathcal{K}\bigr{)}=(1+\delta)^{d}\operatorname{vol}(\mathcal{K})\leq e^{t}\operatorname{vol}(\mathcal{K})$. Taking $c_{1}^{2}=\frac{\log\log Z}{2\log Z}$, $t=\sqrt{8}\log\log Z$, and $N=Z(\log Z)^{4}$, we can bound the last line by $\frac{\eta}{mM}$. Therefore,
@@
We now bound the expected number of trials per iteration. Let $X$ be the minimum of the threshold $N$ and the number of trials until the first success. Then the expected number of trials per step is bounded by $M\mathbb{E}_{\pi^{Y}}[X]$ due to $\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$. Thus,
@@
$\leq e^{t}+N\exp(-\Omega(t^{2}))\leq(\log Z)^{3}+3(\log Z)^{4}=\mathcal{O}\Bigl{(}\bigl{(}\log\frac{mM}{\eta}\bigr{)}^{4}\Bigr{)}\,.$
Therefore, the expected number of trials per step is $\mathcal{O}\bigl{(}M(\log\frac{mM}{\eta})^{4}\bigr{)}$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎

### v2

@@ 3.2 Contraction and mixing
$\mu_{h-\epsilon}^{\leftarrow}(x)=\tilde{\mu}_{h-\epsilon}^{\leftarrow}(x)=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(x)}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
This follows from that if we consider system started at time $\epsilon>0$, with initial distribution $\mu_{\epsilon}$, then we obtain the above through the Bayesian perspective on the forward and reverse heat semigroups, elaborated in §D.
We note that $\tilde{\mu}_{h}^{\leftarrow}$ defined below is obviously integrable with respect to $x$, since this is a probability density function of $\mu_{h}Q$, where $Q$ is the backward kernel:
$\tilde{\mu}_{h}^{\leftarrow}(x):=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}\nu(x)}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
Hence, $\tilde{\mu}_{h}^{\leftarrow}$ is well-defined over the support of $\nu$ almost surely.
We now show the pointwise convergence of $\tilde{\mu}_{h-\epsilon}^{\leftarrow}$ to $\tilde{\mu}_{h}^{\leftarrow}$ as $\epsilon\to 0$. Note that $\nu P_{\epsilon}\to\nu$, as $\nu$ has a bounded support. Also, the denominator is independent of $\epsilon$ due to
$\frac{1}{(2\pi(h-\epsilon))^{d/2}}\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z=\mathcal{N}\bigl{(}0,(h-\epsilon)I\bigr{)}*\nu P_{\epsilon}=\nu*\mathcal{N}(0,hI)\,.$
@@
$\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}$ $\leq\Bigl{(}\frac{h}{h-\epsilon}\Bigr{)}^{d/2}\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}$
$\lesssim\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,.$
Since the last bound is integrable with respect to $\mu_{h}$, the dominated convergence theorem implies
$\lim_{\epsilon\to 0}\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,,$
Thus, the pointwise convergence follows. Note that if we take $\nu(x)=\pi^{X}(x)=\frac{\mathds{1}_{\mathcal{K}}(x)}{\operatorname{vol}(\mathcal{K})}$, then $\tilde{\mu}_{h}^{\leftarrow}$ is the distribution of the backward step of our algorithm. In particular, this corresponds to first sampling $x\sim\mu_{h}$, then $y\sim Q_{h}^{\nu,h}(\cdot|x)$, which is precisely the law of $\mu_{h}^{\leftarrow}$ given by ($\mathsf{FP\text{-}BH}$).
As for the second statement, it follows from Scheffé’s lemma [Bil95, Theorem 16.12] that the pointwise convergence of $\mu_{h-\varepsilon}^{\leftarrow}\to\mu_{h}^{\leftarrow}$ leads to its $\mathsf{TV}$-convergence, which in turn implies the weak convergence. It follows from lower semicontinuity of $D_{f}$ [AFP00, Theorem 2.34] that the weak convergence ensures $D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu^{\leftarrow}_{h-t})$. ∎
@@ 3.3 Failure probability and wasted steps
We begin by defining a suitable version of local conductance [KLS97].
@@
$\ell(x)\stackrel{{\scriptstyle\mathrm{{\scriptscriptstyle def}}}}{{=}}\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\lVert x-y\rVert^{2})\,\mathrm{d}y}{\int_{\mathbb{R}^{d}}\exp(-\frac{1}{2h}\lVert x-y\rVert^{2})\,\mathrm{d}y}=\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\lVert x-y\rVert^{2})\,\mathrm{d}y}{(2\pi h)^{d/2}}\,.$
@@
For a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ containing a unit ball $B_{1}(0)$,
$\pi^{Y}(\mathcal{K}_{\delta}^{c})\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+\delta d\bigr{)}\,.$
@@
For $y\in\partial\mathcal{K}_{\delta}$, we can take the supporting half-space $H(y)$ at $\mathsf{proj}_{\mathcal{K}}(y)$ containing $\mathcal{K}$, due to convexity of $\mathcal{K}$. Then,
$\pi^{Y}(\mathcal{K}_{\delta}^{c})$ $=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{\mathcal{K}}\frac{\exp(-\frac{1}{2h}\lVert y-x\rVert^{2})}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y$
$\leq\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{H(y)}\frac{\exp(-\frac{1}{2h}\lVert y-x\rVert^{2})}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y$
$=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{z^{2}}{2h})}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y\,.$ (3.3)
@@
$\mathsf{T}(s):=\mathbb{P}_{\mathcal{N}(0,h)}(Z\geq s)=1-\Phi(h^{-1/2}s)\,,$
where $\Phi$ is the CDF of the standard Gaussian. By the co-area formula and integration by parts,
$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{1}{2h}z^{2})}{\sqrt{2\pi h}}\,\mathrm{d}z\mathrm{d}y=\int_{\delta}^{\infty}\mathsf{T}(s)\operatorname{vol}(\partial\mathcal{K}_{s})\,\mathrm{d}s$
@@
Recall that $\mathsf{T}(s)\leq\frac{1}{2}\exp(-\frac{1}{2}(h^{-1/2}s)^{2})$ for $h^{-1/2}s\geq 0$ due to a standard tail bound on a Gaussian distribution. This tail bound, combined with
@@
$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{1}{2h}z^{2})}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y$ $\leq\frac{1}{\sqrt{2\pi h}}\int_{\delta}^{\infty}\exp\bigl{(}-\frac{s^{2}}{2h}\bigr{)}\bigl{(}\underbrace{(1+s)^{d}}_{\leq\exp(sd)}-1\bigr{)}\operatorname{vol}(\mathcal{K})\,\mathrm{d}s$
@@
$\underset{(i)}{\leq}\operatorname{vol}(\mathcal{K})\exp(hd^{2}/2)\exp\bigl{(}-\frac{(\delta-hd)^{2}}{2h}\bigr{)}$
$=\operatorname{vol}(\mathcal{K})\exp\bigl{(}-\frac{\delta^{2}}{2h}+\delta d\bigr{)}\,,$
where in $(i)$ we used the tail bound for a Gaussian. ∎
This core lemma suggests taking $\delta=\nicefrac{{t}}{{d}}$ and $h=\nicefrac{{c}}{{d^{2}}}$ for some $t,c>0$, under which we have
$\pi^{Y}(\mathcal{K}_{\delta}^{c})\leq\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}\,.$
@@
$\ell(y)\leq\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{1}{2h}z^{2})}{\sqrt{2\pi h}}\,\mathrm{d}z=\mathbb{P}_{\mathcal{N}(0,h)}(Z\geq\delta)\leq\exp(-\Omega(t^{2}))\,.$
@@
$\leq\pi^{Y}(\mathcal{K}_{\delta}^{c})+\int_{[\ell\geq N^{-1}\log(3mM/\eta)]}\exp(-\ell N)\,\mathrm{d}\pi^{Y}$
$\qquad\qquad+\int_{\mathcal{K}_{\delta}\cap[\ell<N^{-1}\log(3mM/\eta)]}\frac{\ell(y)}{\operatorname{vol}(\mathcal{K})}\,\mathrm{d}y$
$\leq\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}+\frac{\eta}{3mM}+\frac{\log(3mM/\eta)}{N}\,\frac{\operatorname{vol}(\mathcal{K}_{\delta})}{\operatorname{vol}(\mathcal{K})}$
$\leq\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}+\frac{\eta}{3mM}+\frac{e^{t}}{N}\,\log\frac{3mM}{\eta}\,,$
where we used $\operatorname{vol}(\mathcal{K}_{\delta})\subset\operatorname{vol}\bigl{(}(1+\delta)\mathcal{K}\bigr{)}=(1+\delta)^{d}\operatorname{vol}(\mathcal{K})\leq e^{t}\operatorname{vol}(\mathcal{K})$. Taking $c=\frac{\log\log Z}{2\log Z}$, $t=\sqrt{8}\log\log Z$, and $N=Z(\log Z)^{4}$, we can bound the last line by $\frac{\eta}{mM}$. Therefore,
@@
We now bound the expected number of trials per iteration. Let $X$ be the minimum of the threshold $N$ and the number of trials until the first success. Then the expected number of trials per step is bounded by $M\mathbb{E}_{\pi^{Y}}X$ since $\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$. Thus,
@@
$\leq e^{t}+N\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}\leq(\log Z)^{3}+3(\log Z)^{4}=\mathcal{O}\bigl{(}\log^{4}\frac{mM}{\eta}\bigr{)}\,.$
Therefore, the expected number of trials per step is $\mathcal{O}(M\log^{4}\frac{mM}{\eta})$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎